OpenStudy (anonymous):

NEED HELP find the Exponential equation of the form f(x)=c*a^x that goes thru the points (2,12) and (5,32)

OpenStudy (anonymous):

it looks like \[f(x)=c\times a^x\] you need \(c\) and \(a\)

OpenStudy (anonymous):

you know that \[f(2)=c\times a^2=12\] and \[f(5)=c\times a^5=32\] lets see if we can find \(a\) first

OpenStudy (anonymous):

if you diivide you get \[\frac{c\times a^5}{c\times a^2}=\frac{32}{12}\]or \[a^3=\frac{8}{3}\]

OpenStudy (anonymous):

that makes \[a=\frac{2}{\sqrt[3]{3}}\]